# A unified harmonic framework for dark siren cosmology

### 13 Mar 2026 v1

@@ I Introduction
One of the primary limitations of the galaxy catalog method is the incompleteness of galaxy catalogs at moderate to high redshifts [118, 63, 35, 64]. Because the galaxy catalog method typically assumes a uniform distribution for the unobserved galaxies, little to no information is gained on the redshift of the GW source at these distances. This ignores the clustering of galaxies, as in principle the observed galaxy field can be used to inform the field of missing galaxies. On the other hand, techniques which directly infer cosmology from the cross-correlation statistics of GWs and galaxies have been proposed as a complementary method to dark siren techniques. These cross-correlation methods generally aim to constrain cosmology by cross-correlating GWs in luminosity distance space with galaxies in redshift space, most straightforwardly via the 3x2pt analysis of the $GW$, GW–galaxy, and galaxy–galaxy power spectra [97, 102]. Alternate methods have also been proposed, including a 6x2pt extension that includes weak gravitational lensing [15], inference with the GW–galaxy power spectra conditioned on the observed galaxy overdensity field [93, 31, 92], and using 2-point correlations to predict local galaxy overdensities near observed GWs [61]. There is also significant interest in the GW clustering itself, both to inform the large-scale structure (LSS) of the universe as well as to infer the GW tracer bias, which could provide insights on the host environments and formation origins of GW sources [94, 138, 133, 80, 73, 39].
@@ III.2.1 The 3D power spectrum
Substituting the Poisson variance $\langle s^{X}(\mathbf{r})s^{X}(\mathbf{r}^{\prime})\rangle=\frac{\delta^{(D)}(\mathbf{r}-\mathbf{r}^{\prime})}{\langle n^{X}(\mathbf{r})\rangle}$ and the definition of the power spectrum in Equation 14 yields
@@ Appendix A Recovering the standard galaxy catalog method from a GRF
$\begin{split}&p(\mathbf{a}^{G}[\bar{z}]|\{z\},\{z^{g,\text{tot}}\},\mathbf{a}^{g,\text{tot}}[\bar{z}],\mathbf{\Lambda})=\\ &\qquad\prod_{\ell,m}p(a_{\ell m}^{G}[\bar{z}]|n^{G}[\bar{z}],n^{g,\text{tot}}[\bar{z}],a_{\ell m}^{g,\text{tot}}[\bar{z}],\mathbf{\Lambda})\end{split}$ (107)
@@
where $\mathbf{a}$ refers to the vector of $a_{\ell m}$ over $(\ell,m)$, taking advantage of the separation of harmonic modes. We model the GW field as a Poisson realization of the total galaxy field
$a_{\ell m}^{G}[\bar{z}]=a_{\ell m}^{g,\text{tot}}[\bar{z}]+s_{\ell}^{G}[\bar{z}]$ (108)
where $s_{\ell}^{G}[\bar{z}]=1/\langle n^{G}[\bar{z}]\rangle$ is the shot noise of this shell.
The conditional distribution of a multivariate Gaussian is also a multivariate Gaussian, with a modified mean and variance. To compute what those are, we need the quantities
@@
$\begin{split}C_{\ell}^{G\,g,\mathrm{tot}}[\bar{z}]&=\left\langle a_{\ell m}^{G}[\bar{z}]^{*}a_{\ell m}^{g,\text{tot}}[\bar{z}]\right\rangle\\ &=\left\langle\left(a_{\ell m}^{g,\text{tot}}[\bar{z}]+s_{\ell}^{G}[\bar{z}]\right)^{*}a_{\ell m}^{g,\text{tot}}[\bar{z}]\right\rangle\\ &=\bar{C}_{\ell}^{gg}[\bar{z}]+\frac{1}{\langle n^{g,\text{tot}}[\bar{z}]\rangle}\end{split}$ (109)
and (via a similar calculation)
$C_{\ell}^{GG}[\bar{z}]=\bar{C}_{\ell}^{gg}[\bar{z}]+\frac{1}{\langle n^{G}[\bar{z}]\rangle}$ (110)
@@
Equation 109 was derived modeling GWs as a Poisson point process of the total galaxy field. Equation 109 is what we ought to expect; GWs have the same clustering patterns as the galaxies they are drawn from, and we can recognize the cross shot noise term in the last line (as in Equation 38). Equation 110 is also what we expect from a Poisson point process which traces the galaxies. Note that here, $\overline{C}_{\ell}^{gg}=\overline{C}_{\ell}^{Gg}=\overline{C}_{\ell}^{GG}$ because we have assumed that GWs are uniformly drawn from the galaxies, and that observed and unobserved galaxies do not cluster differently. A more flexible model is easily accommodated by modeling a luminosity-dependent galaxy bias along with selection effects, or more simply by allowing GWs and galaxies to have separate biases.
Using Equations 109 and 110, the distribution of the GW overdensity field conditioned on the observed galaxy overdensity field is
@@
$\Sigma_{\ell m}$ $=\frac{1}{\langle n^{G}[\bar{z}]\rangle}$ (111c)
Notice that the variance of the GW field is now a scale-independent shot noise; we’ve showed explicitly that conditioning on a fixed background of galaxies removes 2-point correlations, as we discussed above! Transforming Equation 111 to real space, one can show that the probability of a GW being found in a given infinitesimal shell $\bar{z}$ and pixel $\bar{\bm{\Omega}}$ is given by
@@
As discussed in the main text, the correct thing to do would be to condition only on the observed galaxies $a_{\ell m}^{g}$, rather than the entire galaxy field $a_{\ell m}^{g,\text{tot}}$. To derive this, we need to compute $C_{\ell}^{Gg}[\bar{z}]$, which can be derived as follows:
$\begin{split}C_{\ell}^{Gg}[\bar{z}]&=\left\langle a_{\ell m}^{G}[\bar{z}]^{*}a_{\ell m}^{g}[\bar{z}]\right\rangle=\left\langle\left(\bar{a}_{\ell m}^{G}[\bar{z}]+s_{\ell}^{G}[\bar{z}]\right)^{*}a_{\ell m}^{g}[\bar{z}]\right\rangle\\ &=p(D^{g}|\bar{z})\left(\bar{C}_{\ell}^{gg}[\bar{z}]+\frac{1}{n^{g}}\right)+\big(1-p(D^{g}|\bar{z})\big)\bar{C}_{\ell}^{gg}[\bar{z}]\\ &=\bar{C}_{\ell}^{gg}[\bar{z}]+\frac{p(D^{g}|\bar{z})}{n^{g}[\bar{z}]}.\end{split}$ (113)
This follows from Equation 56; the second term in the last line can be identified as the cross shot noise (as in Equation 38)
Using Equations 110 and 113, the distribution of the GW field conditioned only on the observed galaxies is
$\mathcal{L}(a_{\ell m}^{G}|$ $\,\mathbf{\Lambda},n^{G},n^{g},a_{\ell m}^{g})=\mathcal{N}\left[a_{\ell m}^{G};\mu_{\ell m},\Sigma_{\ell m}\right]$ (114a)
$\mu_{\ell m}$ $=\frac{\overline{C}_{\ell}^{gg}+\frac{p(D^{g}|z,\mathbf{\Lambda})}{n^{g}}}{\overline{C}_{\ell}^{gg}+\frac{1}{n^{g}}}a_{\ell m}^{g}$ (114b)
$\Sigma_{\ell m}$ $=\overline{C}_{\ell}^{gg}+\frac{1}{n^{G}}-\frac{\left[\overline{C}_{\ell}^{gg}+\frac{p(D^{g}|z,\mathbf{\Lambda})}{n^{g}}\right]^{2}}{\overline{C}_{\ell}^{gg}+\frac{1}{n^{g}}}.$ (114c)
With the usual assumption that $N^{g}\gg N^{G}$, this reduces to Equation 111 only in the case of galaxy catalog completeness, i.e., $p(D^{g}|z,\mathbf{\Lambda})=1$. If all possible host galaxies are known, then 2-point correlations do not provide any new information for the possible GW host locations.
@@ Appendix B Beam window functions
For a given realization of the universe, the probability density map under noise of the MAP position of some observed GW $i$ is given by $p^{(i)}(\bm{\Omega})$, which we model as an isotropic Gaussian of width $\sigma_{i}$ centered at its true location $\bm{\Omega}_{i}$.
@@
Let $d^{(i)}_{\ell m}$ be the SHT of $p^{(i)}(\bm{\Omega})$. The true map of this event and its spherical transform are given by
$\rho_{i}(\bm{\Omega})$ $=\delta^{(D)}(\bm{\Omega}-\bm{\Omega}_{i})$ (116)
@@
and thus $p^{(i)}(\bm{\Omega})$ is written in harmonic space as
@@
This expression simplifies if $p(\bm{\Omega}|\bm{\Omega}_{i})$ is isotropic and spatially invariant, i.e., it can be written as a function of $\cos\theta=\bm{\Omega}\cdot\bm{\Omega}_{i}$. We can do this by expanding $P(\cos\theta)$ in the complete basis of the Legendre polynomials $\mathcal{P}_{\ell}(\cos\theta)$
@@
In particular, a Gaussian and isotropic localization error with width $\sigma$
$P(\cos\theta)\propto\exp\left[-\frac{(\cos\theta-1)^{2}}{2\sigma}\right]$ (122)
has a corresponding expansion
$B_{\ell}(\sigma)=\exp\left[-\frac{1}{2}\ell(\ell+1)\sigma^{2}\right]$ (123)
which yields the expression in Equation 68. Although we limit ourselves to isotropic and Gaussian localization errors in Section V.1, our formalism can be extended to more general localization errors using Equation 118.

### 16 Apr 2026 v2

@@ I Introduction
One of the primary limitations of the galaxy catalog method is the incompleteness of galaxy catalogs at moderate to high redshifts [120, 63, 35, 64, 106]. Because the galaxy catalog method typically assumes a uniform distribution for the unobserved galaxies, little to no information is gained on the redshift of the GW source at these distances. This ignores the clustering of galaxies, as in principle the observed galaxy field can be used to inform the field of missing galaxies. On the other hand, techniques which directly infer cosmology from the cross-correlation statistics of GWs and galaxies have been proposed as a complementary method to dark siren techniques. These cross-correlation methods generally aim to constrain cosmology by cross-correlating GWs in luminosity distance space with galaxies in redshift space, most straightforwardly via the 3x2pt analysis of the $GW$, GW–galaxy, and galaxy–galaxy power spectra [98, 103]. Alternate methods have also been proposed, including a 6x2pt extension that includes weak gravitational lensing [15], inference with the GW–galaxy power spectra conditioned on the observed galaxy overdensity field [94, 31, 93], and using 2-point correlations to predict local galaxy overdensities near observed GWs [61]. There is also significant interest in the GW clustering itself, both to inform the large-scale structure (LSS) of the universe as well as to infer the GW tracer bias, which could provide insights on the host environments and formation origins of GW sources [95, 141, 135, 81, 74, 39].
@@ III.2.1 The 3D power spectrum
where $s^{X}(\mathbf{r})$ is defined as the difference between the expected and realized overdensity fields. Assuming homogeneity and isotropy and substituting the Poisson variance $\langle s^{X}(\mathbf{r})s^{X}(\mathbf{r}^{\prime})\rangle=\frac{\delta^{(D)}(\mathbf{r}-\mathbf{r}^{\prime})}{\langle n^{X}(\mathbf{r})\rangle}$ and the definition of the power spectrum in Equation 14 yields [^2] Here, we assume that the shot noise is decoupled from the expected tracer field $\overline{\delta}^{X}(\mathbf{r})$, which is true in the Gaussian limit of Poisson noise ($\langle n^{X}\rangle\gg 1$). Thus, we can average the shot noise over realizations of the field as we have done above (see e.g. [72]).
@@ Appendix A Recovering the standard galaxy catalog method from a GRF
$\begin{split}&p\left(\left\{a_{\ell m}^{G}[\bar{z}]\right\}\Big|\{z\},\{z^{g,\text{tot}}\},\left\{a_{\ell m}^{g,\text{tot}}[\bar{z}]\right\},\mathbf{\Lambda}\right)=\\ &\qquad\prod_{\ell,m}p(a_{\ell m}^{G}[\bar{z}]|n^{G}[\bar{z}],n^{g,\text{tot}}[\bar{z}],a_{\ell m}^{g,\text{tot}}[\bar{z}],\mathbf{\Lambda})\end{split}$ (107)
@@
where $\{a_{\ell m}\}$ refers to the set of $a_{\ell m}$ over $(\ell,m)$, taking advantage of the separation of harmonic modes. We model the GW field as a Poisson realization of the total galaxy field
$a_{\ell m}^{G}[\bar{z}]=a_{\ell m}^{g,\text{tot}}[\bar{z}]+s_{\ell m}^{G}[\bar{z}]$ (108)
where $s_{\ell m}^{G}[\bar{z}]$ is the shot noise of GWs being drawn from the galaxies of this shell, which averaged over all possible realizations of the universe is a scale-independent shot noise
$\left\langle\left(s_{\ell m}^{G}[\bar{z}]\right)^{2}\right\rangle=1/\langle n^{G}[\bar{z}]\rangle.$ (109)
Both $a_{\ell m}^{G}$ and $a_{\ell m}^{g,\text{tot}}$ refer to realized tracer fields, and thus $a_{\ell m}^{g,\text{tot}}$ itself also implicitly contains shot noise, i.e.
$\left\langle\left|a_{\ell m}^{g,\text{tot}}\right|^{2}\right\rangle=C_{\ell}^{gg}=\overline{C}_{\ell}^{gg}[\bar{z}]+\frac{1}{\langle n^{g,\text{tot}}[\bar{z}]\rangle}.$ (110)
To compute the distribution of the GW overdensity field conditioned on the galaxy overdensity field, we can use the standard formulae for the conditional distribution of a multivariate Gaussian. For jointly distributed vectors $\mathbf{z}\equiv(\mathbf{x},\mathbf{y})$ with mean $\langle\mathbf{z}\rangle\equiv\left(\langle\mathbf{x}\rangle,\langle\mathbf{y}\rangle\right)$ and covariance
@@
$\left\langle\mathbf{z}^{\dagger}\mathbf{z}\right\rangle=\begin{pmatrix}\mathbf{\Sigma}_{\mathrm{xx}}&\mathbf{\Sigma}_{\mathrm{xy}}\\ \mathbf{\Sigma}_{\mathrm{xy}}^{\dagger}&\mathbf{\Sigma}_{\mathrm{yy}}\end{pmatrix}\equiv\begin{pmatrix}\left\langle\mathbf{x}^{\dagger}\mathbf{x}\right\rangle&\left\langle\mathbf{x}^{\dagger}\mathbf{y}\right\rangle\\ \left\langle\mathbf{y}^{\dagger}\mathbf{x}\right\rangle&\left\langle\mathbf{y}^{\dagger}\mathbf{y}\right\rangle\end{pmatrix}$, (111)
the distribution of $\mathbf{x}$ conditioned on $\mathbf{y}$ is a multivariate Gaussian with a modified mean and covariance matrix
$\begin{split}\langle\mathbf{x}\rangle_{\mathbf{y}}&=\langle\mathbf{x}\rangle+\mathbf{\Sigma}_{\mathrm{xy}}\mathbf{\Sigma}_{\mathrm{yy}}^{-1}\left(\mathbf{y}-\langle\mathbf{y}\rangle\right)\\ \langle\mathbf{x}^{\dagger}\mathbf{x}\rangle_{\mathbf{y}}&=\mathbf{\Sigma}_{\mathrm{xx}}-\mathbf{\Sigma}_{\mathrm{xy}}\mathbf{\Sigma}_{\mathrm{yy}}^{-1}\mathbf{\Sigma}_{\mathrm{xy}}^{\dagger}.\end{split}$ (112)
Hence, if we take $\mathbf{x}$ to be our GW field $\{a_{\ell m}^{G}[\bar{z}]\}$ and $\mathbf{y}$ to be our galaxy field $\{a_{\ell m}^{g,\text{tot}}[\bar{z}]\}$ at redshift $\bar{z}$, we need the theoretical power spectra $C_{\ell}^{G\,g,\mathrm{tot}}[\bar{z}]$ and $C_{\ell}^{GG}[\bar{z}]$ in order to compute Equation 107. Using Equation 108, we have
$\begin{split}C_{\ell}^{G\,g,\mathrm{tot}}[\bar{z}]&=\left\langle a_{\ell m}^{G}[\bar{z}]^{*}a_{\ell m}^{g,\text{tot}}[\bar{z}]\right\rangle\\ &=\left\langle\left(a_{\ell m}^{g,\text{tot}}[\bar{z}]+s_{\ell m}^{G}[\bar{z}]\right)^{*}a_{\ell m}^{g,\text{tot}}[\bar{z}]\right\rangle\\ &=C_{\ell}^{gg}[\bar{z}]=\overline{C}_{\ell}^{gg}[\bar{z}]+\frac{1}{\langle n^{g,\text{tot}}[\bar{z}]\rangle}\end{split}$ (113)
where we have used the standard approximation that the shot noise is decoupled from the underlying field from which it is drawn $\langle s_{\ell m}^{G}a_{\ell m}^{g,\text{tot}}\rangle=0$, valid for $\langle n^{G}\rangle\gg 1$. A similar calculation yields
$C_{\ell}^{GG}[\bar{z}]=\overline{C}_{\ell}^{gg}[\bar{z}]+\frac{1}{\langle n^{G}[\bar{z}]\rangle}$ (114)
Note that these theoretical power spectra are constructed by definition from the expectation value of the fields over all realizations of the universe; we condition upon a realization of the galaxy field later using Equation 112.
@@
Equation 113 is what we ought to expect; GWs have the same clustering patterns as the galaxies they are drawn from, and we can recognize the cross shot noise term in the last line (as in Equation 38). Equation 114 is also what we expect from a Poisson point process which traces the galaxies. Furthermore, $\overline{C}_{\ell}^{gg}=\overline{C}_{\ell}^{Gg}=\overline{C}_{\ell}^{GG}$ because we have assumed that GWs are uniformly drawn from the galaxies, and that observed and unobserved galaxies do not cluster differently. A more flexible model is easily accommodated by modeling a luminosity-dependent galaxy bias along with selection effects, or more simply by allowing GWs and galaxies to have separate biases.
Equation 43 gives us a joint multivariate normal distribution on the GW and galaxy overdensity fields. Then, substituting Equations 113 and 114 into Equation 112 yields our desired conditional distribution
@@
$\sigma_{\ell m}=\frac{1}{\langle n^{G}[\bar{z}]\rangle}-\frac{1}{\langle n^{g}[\bar{z}]\rangle}$ $\approx\frac{1}{\langle n^{G}[\bar{z}]\rangle}$ (115c)
assuming $n^{G}\gg n^{g}$. Equation 115 is particularly simple because $C_{\ell}^{G\,g,\mathrm{tot}}=C_{\ell}^{gg}$. Notice that the variance of the GW field is now simply a scale-independent shot noise. Thus, we have now showed explicitly that conditioning on a fixed background of galaxies removes 2-point correlations, as discussed in Section IV.2.
Transforming Equation 115 to real space, one can show that the probability of a GW being found in a given infinitesimal shell $\bar{z}$ and pixel $\bar{\bm{\Omega}}$ is given by
@@
As discussed in the main text, the correct thing to do would be to condition only on the observed galaxies $\{a_{\ell m}^{g}[\bar{z}]\}$, rather than the entire galaxy field $\{a_{\ell m}^{g,\text{tot}}[\bar{z}]\}$. To compute this, we need the quantity $C_{\ell}^{Gg}[\bar{z}]$. Substituting Equation 56, we have
$\begin{split}C_{\ell}^{Gg}[\bar{z}]&=\left\langle a_{\ell m}^{G}[\bar{z}]^{*}a_{\ell m}^{g}[\bar{z}]\right\rangle=\left\langle\left(a_{\ell m}^{g,\text{tot}}[\bar{z}]+s_{\ell m}^{G}[\bar{z}]\right)^{*}a_{\ell m}^{g}[\bar{z}]\right\rangle\\ &=p(D^{g}|\bar{z})C_{\ell}^{gg}[\bar{z}]+\big(1-p(D^{g}|\bar{z})\big)C_{\ell}^{g\,g,\mathrm{miss}}[\bar{z}]\\ &=p(D^{g}|\bar{z})\left(\overline{C}_{\ell}^{gg}[\bar{z}]+\frac{1}{n^{g}}\right)+\big(1-p(D^{g}|\bar{z})\big)\overline{C}_{\ell}^{gg}[\bar{z}]\\ &=\overline{C}_{\ell}^{gg}[\bar{z}]+\frac{p(D^{g}|\bar{z})}{n^{g}[\bar{z}]}.\end{split}$ (117)
The second term in the last line can again be identified as the cross shot noise in Equation 38. As above, we assume for simplicity that the missing and galaxies cluster with the same power spectrum $\overline{C}_{\ell}^{gg}$, which could be generalized with respective bias parameters.
Using Equations 114 and 117, the distribution of the GW overdensity field conditioned only on the observed galaxies is
$p(a_{\ell m}^{G}[\bar{z}]|$ $\,n^{G}[\bar{z}],n^{g}[\bar{z}],a_{\ell m}^{g}[\bar{z}],\mathbf{\Lambda})=\mathcal{N}\left[a_{\ell m}^{G};\mu_{\ell m},\sigma_{\ell m}\right]$ (118a)
$\mu_{\ell m}$ $=\frac{\overline{C}_{\ell}^{gg}+\frac{p(D^{g}|z,\mathbf{\Lambda})}{\langle n^{g}[\bar{z}]\rangle}}{\overline{C}_{\ell}^{gg}+\frac{1}{\langle n^{g}[\bar{z}]\rangle}}a_{\ell m}^{g}$ (118b)
$\sigma_{\ell m}$ $=\overline{C}_{\ell}^{gg}+\frac{1}{\langle n^{G}[\bar{z}]\rangle}-\frac{\left[\overline{C}_{\ell}^{gg}+\frac{p(D^{g}|z,\mathbf{\Lambda})}{\langle n^{g}[\bar{z}]\rangle}\right]^{2}}{\overline{C}_{\ell}^{gg}+\frac{1}{\langle n^{g}[\bar{z}]\rangle}}.$ (118c)
With the usual assumption that $n^{g}\gg n^{G}$, this reduces to Equation 115 only in the case of galaxy catalog completeness, i.e., $p(D^{g}|z,\mathbf{\Lambda})=1$. Thus, if all possible host galaxies are known, then 2-point correlations do not provide any new information for the possible GW host locations.
@@ Appendix B Beam window functions
For a given realization of the universe, the probability density map under noise of the MAP position of some observed GW $i$ is given by $p^{(i)}(\bm{\Omega})$, which we model as an isotropic Gaussian of width $\sigma_{i}$ centered at its true location $\bm{\Omega}_{i}$. Thus $p^{(i)}(\bm{\Omega})$ is a function of $\bm{\Omega},\bm{\Omega}_{i}$, and to make this manifest we will now use $p(\bm{\Omega}|\bm{\Omega}_{i})=p^{(i)}(\bm{\Omega})$ for the probability distribution of the observed position of the $i$th event.
@@
Let $d^{(i)}_{\ell m}$ be the SHT of $p(\bm{\Omega}|\bm{\Omega}_{i})$. The true number density field of this event and its spherical transform are given by
$n_{i}(\bm{\Omega})$ $=\delta^{(D)}(\bm{\Omega}-\bm{\Omega}_{i})$ (120)
@@
and thus $p(\bm{\Omega}|\bm{\Omega}_{i})$ is written in harmonic space as
@@
This expression simplifies if $p(\bm{\Omega}|\bm{\Omega}_{i})$ is isotropic and spatially invariant, i.e., it can be written as a function of $\cos\theta=\bm{\Omega}\cdot\bm{\Omega}_{i}$. We can do this by expanding $p(\bm{\Omega}|\bm{\Omega}_{i})\equiv p(\cos\theta)$ in the complete basis of the Legendre polynomials $\mathcal{P}_{\ell}(\cos\theta)$
@@
In particular, a Gaussian and isotropic localization error with RMS width $\sigma$ can be written as
$p(\cos\theta)\propto\exp\left[-\frac{1-\cos\theta}{\sigma^{2}}\right]\approx\exp\left[-\frac{\theta^{2}}{2\sigma^{2}}\right]$ (126)
where the small angle approximation is valid for $\sigma\lesssim 0.1$. This has a corresponding expansion (to leading order in $\ell^{2}\sigma^{2}$) [137]
$B_{\ell}\approx\exp\left[-\frac{1}{2}\ell(\ell+1)\sigma^{2}\right].$ (127)
valid for $\ell^{2}\sigma^{2}\ll 1$. Becauase the signal is suppressed for $\ell^{2}\sigma^{2}\gtrsim 1$ anyway, higher-order corrections are not needed. This yields the expression in Equation 68. Although we limit ourselves to isotropic and Gaussian localization errors in Section V.1, our formalism can be extended to more general localization errors using Equation 122.